### v1

# Infinite-Horizon Optimal Wireless Control Over Shared State-Dependent Fading Channels for IIoT Systems

Abstract
Heterogeneous systems consisting of a multiloop wireless control system (WCS) and a mobile agent system (MAS) are ubiquitous in Industrial Internet of Things systems. Within these systems, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance. This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to state and input constraints. Firstly, the state-dependent fading channel is modeled, which characterizes the interference among transmission links, and shows that the probability of a successful transmission for WCS depends on the state of MAS. A necessary and sufficient condition in terms of constrained set stabilization is then established to ensure the Lyapunov-like performance of WCS with expected decay rate. Secondly, using the semi-tensor product of matrices and constrained reachable sets, a criterion is presented to check the constrained set stabilization of MAS and to ensure the performance of WCS. In addition, a constrained optimal state transition graph is constructed to address state and input constraints, by resorting to which the feasibility analysis of the optimal control problem is presented. Finally, an algorithm is proposed for the construction of optimal input sequences by minimum-mean cycles for weighted graph. An illustrative example is provided to demonstrate effectiveness of the proposed method.
keywords:
Constrained set stabilization; heterogeneous Industrial Internet of Things system; infinite-horizon optimal control; shared wireless medium.
thanks: The work was supported by National Key R&D Program of China under the grant 2022YFB3303900 and the NSF of China under the grants 61933009, 62103272. This paper was not presented at any IFAC meeting. Corresponding author S. Zhu.
1 Introduction
1.1 Background and motivation
Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields including smart manufacturing through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011]. These plants consist of multiple physical systems with sensing, computation, communication and actuation capabilities, which flexibly interact with one another and human operators [Ahlen2019, vitturi2019]. Due to such complex interactions among the entities, wired communication will reach its limits. Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming wireless control systems (WCSs). This abundance of wireless devices spawns the need of efficiently sharing the available wireless medium [gatsis2018].
An integral aspect of IIoT in smart manufacturing is the use of mobile agents [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, forming a heterogeneous IIoT system, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019]. In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as collision and dangerous position avoidance. Moreover, positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over, leading to variations in the measured channel gains [Agrawal2014, Quevedo2013]. In order to ensure the performance of WCSs, it is necessary to explicitly examine the influence of mobile agents’ positions on channel states of the WCS, and design constrained agents motion strategy such that the desired performance requirements is guaranteed for the WCS. Such coupling brings heterogeneity to IIoT systems, namely, the WCS is coupled with the mobile agent system (MAS), which is a distinctive feature of IIoT systems compared with traditional WCSs.
Some results considered the dependence of channel state on physical states, which mainly focused on channel characterization, and communication and controller design of WCSs [Agrawal2014, Ahlen2019, Kashiwagi2010]. Few works addressed the heterogeneity feature induced by this dependence. A state-dependent channel model was proposed in [hu2019] to characterize the WCS and MAS coupling, and approximate optimal control and transmission power policies were obtained, under which the WCS achieves both safety and efficiency. This work was generalized in [hu2022] by considering a state-dependent Markov channel model to address time varying data rates, where optimal strategies were obtained by solving a constrained polynomial optimization problem. In these results, the state and input constraints for the MAS were ignored. In addition, only a simplified model consisting of a WCS with a single control loop and a single agent was considered. How to solve the coordination problem of WCSs and MASs in IIoTs with multiple plants being controlled over a shared wireless channel is largely unexplored. The main challenge is to model the fading channel which captures both the interference among transmission links, and the coupling between WCS and MAS, based on which to address the heterogeneity of the whole IIoT system.
1.2 Related work
Since the typical connection density in the IIoT scenario is $10^{6}/km^{2}$, transmission scheduling among sensors over the shared limited wireless medium is an important topic in practice [valerio2021]. For example, static schedules typically specifying periodic communication sequences were proposed to satisfy given control performance requirements such as stability [Hristu2001], controllability and observability [zhang2006] and minimizing linear quadratic objectives [le2011]. Dynamic schedules deciding access to the shared wireless medium dynamically at each step were also extensively studied [donkers2011, gatsis2015, Molin]. For a survey on design and optimization for WCSs, please refer to [Park2018].
As a part of IIoT, WCSs with fading channels have been extensively studied in the past few decades. It is noted that the channel gains used to characterize shadow fading are traditionally modeled as independent identical distributed random processes [Gatsis2014] or Markov chains [zhangq1999], where the network state is assumed to be independent from physical states. With such independency, several results were established for WCSs over fading channels including co-design problem for communication and control systems [Gatsis2014], state estimation [Quevedo2012] and so on. References [gatsis2015] and [gatsis2018] considered shared fading channels, where centralized and distributed channel-aware access mechanisms were respectively presented to meet the Lyapunov-like performance requirements while minimizing the total expected power consumption. It is obvious that these methods on performance analysis of WCS with fading channels fail in the IIoT scenario where the WCS and MAS coupling exists, not to mention the case of constrained MAS.
Few works addressing dependence of network state on physical states have been conducted in IIoT. A state-dependent bursty fading channel model was proposed in [hu2013], based on which distributed switching control [hu2015] and power control schemes [hu2020] were respectively proposed to ensure the performance of vehicular WCSs. However, only homogeneous systems without constraints were considered in these works. In order to model random time variations of wireless environment, [Quevedo2013] introduced a network state process for state estimation, but the network topology was assumed to be fixed. The optimal network topology configuration was then determined for each network state to minimize an expected error covariance measure [Leong2016]. However, the external environment in [Leong2016, Quevedo2013] was modeled as a (semi)-Markov chain which cannot be controlled. Removing the uncontrollability assumptions and modeling the external environment as a Markov decision process, [hu2019] studied the infinite-horizon optimal design problem of communication and control, such that the single-loop WCS with state-dependent channel achieves both safety and efficiency. By formulating this problem as a constrained two-player cooperative game, approximate optimal control and transmission power policies were obtained by solving relaxed convex generalized geometric programs. Similar problem was studied in [hu2022], which considered a generalized state-dependent Markov channel model to address time varying data rates. By solving a constrained polynomial optimization problem, optimal design strategies were obtained. In the above works, safety constraints of the mobile agents were ignored. In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, and coordinate with multiple mobile agents to complete manufacturing tasks, the proposed state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
As for the other part of IIoT, recently, the finite field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014]. The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011]. It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints. Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems. The optimal problem with a set stabilization constraint was considered in [gao2021], where only the finite-horizon optimization until reaching the target set was addressed. Noting that all the above results are established for homogeneous finite-valued systems, i.e., the dynamics of all nodes in the system is the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to performance requirements constraints.
1.3 Contributions
In this paper, we focus on the optimal control of MAS with state and input constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling. A preliminary version of the paper appears in [wangs2022], where only controller design problem is addressed for the MAS to ensure the performance of WCS. The main contributions of the paper are summarized as follows.
First, to take into account the level of shadowing effects on channel gains of the WCS, the finite field network architecture is used to model the operational control for the MAS. The state-dependent fading channel is then modeled to characterize the dependence of wireless transmission of the WCS on mobile agents’ positions, leading to the state-dependent Lyapunov-like performance requirement of the WCS. Based on this model, the infinite-horizon optimal control problem is established for the heterogeneous IIoT system to quantify the tradeoff between WCS’s power consumption and MAS’s control cost, while ensuring the Lyapunov-like performance requirement of the WCS and the safety constraints of the MAS.
Second, optimal input sequences are designed for the MAS to solve the infinite-horizon optimal control problem. A criterion in terms of constrained set stabilization for the MAS is established to ensure the Lyapunov-like performance of WCS, which equivalently converts the optimal control problem into the infinite-horizon optimal constrained set stabilization problem of the MAS. Since each agent updates its states from a finite field, it is difficult to apply classical nonlinear system theory to solve the optimal problem. We adopt the semi-tensor product approach to represent the MAS as a discrete-time bilinear system, based on which a necessary and sufficient condition is presented for the constrained set stabilization of MAS. In order to address state and input constraints of the MAS, a constrained optimal state transition graph is constructed, by analyzing which the feasibility of the considered optimal control problem is proved. Then, an algorithm is proposed to design optimal input sequences by resorting to minimum-mean cycles in weighted graph.
This paper is organized as follows. In Section 2, we characterize the heterogeneous IIoT system and formulate the problem studied in this paper. Section 3 considers Lyapunov-like performance analysis of the WCS. Feasibility analysis and optimal controller design are proceeded in Section 4 and Section 5, respectively. Section 6 demonstrates main results by an illustrative example, which is followed by the conclusion in Section 7. The key notations used in this paper are summarized in Table 1. Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$ [chengqi2011, cheng2011]. Since the semi-tensor product is a generalization of the conventional matrix product, the symbol $\ltimes$” is omitted in most places of this paper when no confusion arises.
Table 1: Notations.
Notations Definitions
$\mathbb{N}_{+}$ Set of natural numbers
$\mathcal{D}_{s}$ Logic domain $\{0,1,\cdots,s-1\}$
$\mathcal{D}_{n}^{m}$ Cartesian product $\underbrace{\mathcal{D}_{n}\times\cdots\times\mathcal{D}_{n}}_{m}$
$I_{s}$ $s$-dimensional identity matrix
$\delta_{s}^{i}$ $i$-th column of $I_{s}$
$[F]_{:,j}$ $j$-th column of matrix $F$
$[F]_{i,j}$ $(i,j)$-th entry of matrix $F$
$\delta_{s}[i_{1}\;i_{2}\;\cdots\;i_{t}]$ Logical matrix $F$ with $[F]_{:,j}=\delta_{s}^{i_{j}}$
${\mathcal{L}}^{n\times m}$ Set of $n\times m$ logical matrices
$\mathbb{R}^{n\times m}$ Set of $n\times m$ real matrices
$Col(F)$ Set $\{[F]_{:,j}:j=1,\cdots,m\}$ for $F\in\mathbb{R}^{n\times m}$
$\ltimes$ Semi-tensor product
$\otimes$ Kronecker product
$|\mathcal{M}|$ Cardinal number of set $\mathcal{M}$
$\mathcal{M}\setminus\mathcal{N}$ Set $\{x\in\mathcal{M}:x\notin\mathcal{N}\}$
2 System model and problem formulation
Figure 1: Framework of a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a factory floor, discretized into $\kappa$ subregions. ‘S’ and ‘A’ respectively represent sensing and actuation capabilities. The local channel state of each control loop $i$ ($\gamma_{i}\in\mathcal{D}_{r}$) is dependent on positions of agents $\alpha_{j}\in\mathcal{D}_{\kappa}$, $j=1,\cdots,n$.
We consider a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a typical smart manufacturing scenario (Fig. 1). In the wireless control architecture, $q$ independent plants are controlled over a shared wireless medium, where the measurements of plant $i$, $i\in\{1,\cdots,q\}$ are wirelessly transmitted to the access point by sensor $i$ to compute the control inputs. At the same time, $n$ mobile agents perform manufacturing tasks in the same workspace. In this process, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over, resulting in the coupling between the WCS and MAS. According to the level of shadowing effects on channel gains of the WCS, partition all possible positions of the mobile agents in the factory floor into $\kappa$ overlapping sets. Each set corresponds to a certain level of shadowing effects. Correspondingly, the whole workspace is discretized into $\kappa$ overlapping planar subregions, called cells $\{0,1,\cdots,\kappa-1\}$.
2.1 Plants
Denote the state of plant $i$ at time $k$ by $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
$x_{i}(k+1)=A_{i}x_{i}(k)+B_{i}v_{i}(k)+\xi_{i}(k),$ (1)
where $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}\times 1}$ is the driving input, and $\{\xi_{i}(k):k\in\mathbb{N}\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
2.2 Sensors and access point
$q$ sensors are placed to monitor the plants. At every slot $k$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not adapting to its local wireless channel state. We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at slot $k$ is $\gamma_{i}(k)\in\mathcal{D}_{r}$. Assume that the wireless communication policy is pre-designed as
$\varphi_{i}(k)=h_{i}(\gamma_{i}(k)),$ (2)
where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(k)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at slot $k$. In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot. The pilot signal may also serve as a synchronization purpose [gatsis2018, hu2011].
When sensor $i$ decides to transmit, it wirelessly transmits measurements of plant $i$ to the access point with a fixed power $\mu_{i}>0$, where the access point plays the role of remote controller responsible for computing control decisions based on the received measurements [ma2018]. Then, control decisions are wirelessly fed back to actuator $i$. These $q$ independent control loops form a multiloop WCS.
2.3 Multiloop WCS
The transmission of measurements and control decisions might fail due to packet decoding errors and packet collisions. In this paper, we focus on the uplink transmission from the sensors to the access point, the idea can be extended into the downlink transmission scenario straightforwardly. Use $\lambda_{i}(k)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time $k$. Then, the measurement $x_{i}(k)$ is used to compute the control input $v_{i}(k)$ by pre-designed control policy at a successful transmission ($\lambda_{i}(k)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$. Otherwise, an open-loop system with system matrix $A_{o,i}$ which may be unstable is obtained. The resulting system can be uniformly described as the following switched model:
$x_{i}(k+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=1,\\ A_{o,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
2.4 Finite-field MAS
Due to the discretization of the workspace, it is the cell that the agents are located in, not the specific position, that directly affects the channel condition of the WCS. In this way, the state of an agent is defined as the index of the cell that the agent is located in, i.e., the state space of each agent is the finite set $\mathcal{D}_{\kappa}$.
At each time step, each agent updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite field model:
$\alpha_{j}(k+1)=\{_{\kappa}\sum\limits_{l\in\bar{\mathcal{I}}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
where ¯$\bar{\mathcal{I}}_{j}=\mathcal{I}_{j}\cup\{j\}$, $\mathcal{I}_{j}$ is the set of in-neighbors of $j$, $\alpha_{j}(k)$, $u_{j}(k)$ respectively denote the state and input of $j$ at time $k$, $a_{j,l}\in\mathcal{D}_{\kappa}$ are the weights, and operations $+_{\kappa}$, ×$\times_{\kappa}$ denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$, respectively [Lidl1996]. In this way, the state $\alpha_{j}(k)$ will evolve over the finite cells $\{0,1,\cdots,\kappa-1\}$. Such finite field architecture has been widely studied to address problems for networks of agents with finite state space [Pasqualetti2014, Sundaram2013]. To ensure production safety such as collision and dangerous position avoidance, we impose state constraints $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}^{n}$ and state-dependent input constraints $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))\subseteq\mathcal{D}_{\kappa}^{n}$. This is one of the differences with the existing results [hu2022, hu2019] which model the mobile agent by an unconstrained Markov decision process. The existence of constraints makes the control of MAS much more complex.
2.5 State-dependent fading channel model
In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and thus influence the local channel state. We call a packet collision happens on system $i$, rendering system $i$’s packet decoding failed, if some other sensor $j\neq i$ transmits in the same time slot. If sensor $i$ transmits and has a collision-free time slot, the success of decoding the packet at the access point depends on the local channel state of link $i$, then the state of the MAS. Such WCS and MAS coupling is modeled by the following state-dependent fading channel:
$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\lambda}_{i}(\alpha),$ (5)
where $\bar{\lambda}_{i}(\alpha):\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$, $\lambda_{i}(k)=1$ indicates the successful transmission of system $i$.
Eq. (5) presents a general state-dependent channel model. Communication policies for sensors and the interference among transmission links can be embedded into this model. In this model, the probability of a successful transmission for WCS (3) depends on the state of MAS (4). It is quite different from the traditional shadow fading model [Gatsis2014, zhangq1999] assuming the channel states independent from the physical states. Such coupling makes the traditional methods on performance analysis of WCSs [gatsis2015, gatsis2018, Hristu2001, le2011, Molin, zhang2006] fail in the scenario in this paper.
2.6 Problem formulation
Considering the state-dependent fading channel model (5), given initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$ of the MAS, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ for MAS (4) such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k\in\mathbb{N}$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
Definition 1.
The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}\subseteq\mathcal{D}_{\kappa}^{n}$ and a positive integer $K$ such that the inequalities
$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]\leq\rho_{i}V_{i}(x_{i}(k))$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+Tr(Q_{i}\Xi_{i}),\forall\ x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ (6)
hold for any $k\geq K$, any $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ and any $i\in\{1,\cdots,q\}$, where $\mathbb{E}$ is the expectation, $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
In (6), $\rho_{i}$ represents desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation. Similar kinds of performance requirements as in (6) have been adopted for WCSs [gatsis2015, gatsis2018]. However, different from them, Definition 1 explicitly explores the state-dependent fading channel.
Our second goal in this paper is to determine optimal input sequence ${\bf{u}}^{\ast}$ such that the average cost function
$J_{{\bf{u}}}(\alpha_{0})=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\sum\limits_{k=0}^{K-1}\Big{\{}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda g(\alpha(k;\alpha_{0},{\bf{u}}),u(k))\Big{\}}$ (7)
reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}^{m+n}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
[…]
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},$
$~{}~{}~{}~{}~{}~{}~{}~{}u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (8)
In the above problem, there are heterogeneous system dynamics constraints (3), (4), WCS’s performance requirements (6) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, which make the resolution of this problem difficult.
Figure 2: Main idea of solving the infinite-horizon optimal control problem (8).
3 Lyapunov-like performance analysis over the state-dependent fading channel
In the optimal control problem (8), Lyapunov-like performance requirements constraints (6) brings about couplings between the WCS and the MAS, making the resolution of problem (8) challenging. To address this issue, we convert constraint (6) for WCS (3) and safety constraints to the constrained set stabilization of MAS (4) using the state-dependent fading channel (5). In this way, all constraints in problem (8) are related to the MAS, which facilitates its optimal input sequence design. Motivated by this observation, the main idea of solving problem (8) is illustrated in Fig. 2.
3.1 Characterization of state-dependent fading channel
In this part, we characterize the state-dependent fading channel in (5) by taking into account of the wireless communication policy and the interference among transmission links.
For fixed wireless communication policy, the probability of successful transmission for link $i$ can be expressed as $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}$, where $\eta_{i}(k)=1$ indicates the successful packet decoding. Due to the discrete channel state, the local channel state in the presence of shadow fading caused by movements of mobile agents is computed as follows:
$\mathbb{P}\{\gamma_{i}(k)=a|\alpha(k)=\alpha\}=\bar{\gamma}_{i}(a,\alpha),i=1,\cdots,q,$ (9)
where $\mathbb{P}$ is the probability, and $\bar{\gamma}_{i}:\mathcal{D}_{r}\times\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$ can be easily obtained by measuring the local fading amplitude for each MAS state according to the statistical methods in [Agrawal2014, Kashiwagi2010] and reference therein. Then, according to (2), we have $\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)$
Since the transmit power is fixed, the probability that the packet of system $i$ is successfully decoded at the access point depends on the local channel state of link $i$, then the state of the MAS as
$\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\eta}_{i}(\alpha),$ (10)
where $\bar{\eta}_{i}:\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$. The interference among transmission links is taken into account by adopting the collision channel model. Then, it holds
$\bar{\eta}_{i}(\alpha)=$$\mathbb{P}\{\hbox{SNR}_{i}(k)\geq\eta_{i}^{0}|\alpha(k)=\alpha,\varphi_{i}(k)=1,$
$\varphi_{j}(k)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(k)=0|\alpha(k)=\alpha\},$
where $\hbox{SNR}_{i}(k)=\mu_{i}\gamma_{i}^{2}(k)/N_{0}$, $N_{0}$ and $\eta_{i}^{0}$ respectively denote the signal-to-noise ratio [hu2019], noise power and the specified decodability threshold of system $i$. Based on the above analysis, we obtain the following result.
Lemma 2.
For the state-dependent fading channel (5), it holds $\bar{\lambda}_{i}(\alpha)=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)\bar{\eta}_{i}(\alpha)$, $\alpha\in\mathcal{C}_{\alpha}$, $i=1,\cdots,q$.
3.2 Lyapunov-like performance analysis of WCS
Based on the characterization of state-dependent fading channel, we can now analyze the Lyapunov-like performance of the WCS. For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ an admissible input sequence in the following.
Definition 3.
Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{M}$ holds for any $k\geq K_{0}$.
Definition 4.
A subset $\mathcal{S}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset (CCIS) of MAS (4), if for any $\alpha(0)=\alpha_{0}\in\mathcal{S}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{S}$ holds for any $k\in\mathbb{N}$.
Obviously, the union of any two CCISs of MAS (4) is another CCIS. The union of all CCISs contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
Theorem 5.
The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}\times 1},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (11)
Proof 3.1.
For each subsystem $i$, since $\mathbb{E}[\xi_{i}(k)]=\mathbf{0}$ and $\lambda_{i}(k)$ is independent of $x_{i}(k)$, by (3), it holds
$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]$
$=$ $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(k)$
+$+\mathbb{P}\{\lambda_{i}(k)=0|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(k)$
$+T_{r}(Q_{i}\Xi_{i}).$
Then, condition (6) subject to the multiloop WCS dynamics (3) is equivalent to the condition that for all $x_{i}(k)\neq\mathbf{0}$, it holds
$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
$\geq$ $\frac{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(k)}{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(k)},$ (12)
which is equivalent to $\bar{\lambda}_{i}(\alpha(k;\alpha_{0},{\bf{u}}))\geq s_{i}$ with $s_{i}$ given in (11), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
Next, we prove that MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$ if and only if it is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. We just need to prove the necessity. This is proved by a contradiction argument. Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence $\bar{{\bf{u}}}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\Omega(s)$.
Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}})\notin I(\Omega(s))$. Since $I(\Omega(s))$ is a largest CCIS, for any $\bar{\alpha}\in\Omega(s)\setminus I(\Omega(s))$ and any admissible input sequence ${\bf{u}}$, there exists an integer $k(\bar{\alpha},{\bf{u}})$ such that $\alpha(k(\bar{\alpha},{\bf{u}});\bar{\alpha},{\bf{u}})\notin\Omega(s)$. Then, it holds
$\alpha(k_{0}+k(\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}}),\tilde{{\bf{u}}});\alpha_{0},\bar{{\bf{u}}})\notin\Omega(s),~{}\tilde{u}(k)=\bar{u}(k+k_{0}),$
which is a contradiction. Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. $\Box$
By (12), $s_{i}$ represents the lower bound of the probability of successful transmission for system $i$ that ensures the desired Lyapunov-like performance given in Definition 1 under the WCS and MAS coupling. In addition, $\Omega(s)$ consists of all constrained states of the MAS satisfying the performance requirements constraint (6) of WCS (3). Thus, for the optimal control problem (8), all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable, that is, problem (8) is equivalently converted into the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
[…]
$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k;\alpha_{0},{\bf{u}})\in I(\Omega(s)),\forall\ k\geq K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},\forall\ k<K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (13)
4 Feasibility analysis
In this section, we first reformulate the MAS in a tractable form, and then present a criterion on the constrained set stabilization of MAS (4). Finally, the feasibility of optimal control problem (13) is analyzed based on the graphical method.
4.1 Algebraic reformulation of the MAS
Due to the highly nonlinear finite field model, it is difficult to apply classical nonlinear system theory to investigate (4). Recently, the semi-tensor product approach has emerged as an elegant tool to deal with such issues [chengqi2011, cheng2011]. The basic idea is to embed the nonlinear finite filed dynamics in (4) into a higher dimensional space where its evolution is linear. In this part, by virtue of the semi-tensor product, we convert MAS (4) into a bilinear equivalent algebraic form, which facilitates the sequential studies.
Set
$\bar{a}_{j,l}=\left\{\begin{array}[]{ll}a_{j,l},&\hbox{if}\ l\in\bar{\mathcal{I}}_{j},\\ 0,&\hbox{otherwise},\end{array}\right.$
and represent $i-1\in\mathcal{D}_{\kappa}$ by the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$. Then, $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in Col(I_{\kappa})$, respectively. Letting $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)$, we have $\mathcal{C}_{\alpha},\mathcal{C}_{u}(\alpha(k))\subseteq Col(I_{N})$, where $N:=\kappa^{n}$. Then, based on similar calculation to [wangs2022], we have the following result.
Lemma 6.
MAS (4) has the following equivalent algebraic form:
$\alpha(k+1)=Fu(k)\alpha(k),$ (14)
where $F\in\mathcal{L}^{N\times N^{2}}$ satisfies $[F]_{:,l}=\ltimes_{j=1}^{n}[F_{j}]_{:,l}$, $l=1,\cdots,N^{2}$, $F_{j}=S_{\kappa}^{n}\ltimes_{l=1}^{n}[I_{\kappa^{l-1}}\otimes(M_{\kappa}\bar{a}_{j,l})]W_{[\kappa,\kappa^{n}]}({\bf{1}}_{\kappa^{j-1}}^{\top}\otimes I_{\kappa}\otimes{\bf{1}}_{\kappa^{n-j}}^{\top})$, $j=1,\cdots,n$, $S_{\kappa}$ and $M_{\kappa}$ respectively represent the structural matrices for $+_{\kappa}$ and ×$\times_{\kappa}$ [Liy2019], ${\bf{1}}_{s}$ represents the $s\times 1$ vector with all entries being $1$, and $W_{[s,t]}$ represents the swap matrix defined as $W_{[s,t]}=[I_{t}\otimes\delta_{s}^{1}~{}\cdots~{}I_{t}\otimes\delta_{s}^{s}]$.
4.2 Criterion of constrained set stabilization
In this part, based on the constrained reachable set, we present a criterion for the constrained set stabilization of MAS (14).
Denote the set of constrained states reachable from $\alpha_{0}$ by $\mathcal{R}(\alpha_{0})$. Then, we have the following result.
Lemma 7.
MAS (14) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$.

### v2

@@
 Heterogeneous systems consisting of a multiloop wireless control system (WCS) and a mobile agent system (MAS) are ubiquitous in Industrial Internet of Things systems.
-Within these systems, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance.
-This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to state and input constraints.
-Firstly, the state-dependent fading channel is modeled, which characterizes the interference among transmission links, and shows that the probability of a successful transmission for WCS depends on the state of MAS.
-A necessary and sufficient condition in terms of constrained set stabilization is then established to ensure the Lyapunov-like performance of WCS with expected decay rate.
-Secondly, using the semi-tensor product of matrices and constrained reachable sets, a criterion is presented to check the constrained set stabilization of MAS and to ensure the performance of WCS.
-In addition, a constrained optimal state transition graph is constructed to address state and input constraints, by resorting to which the feasibility analysis of the optimal control problem is presented.
-Finally, an algorithm is proposed for the construction of optimal input sequences by minimum-mean cycles for weighted graph.
+Within these systems, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance, requiring joint coordination between the WCS and MAS.
+Such coordination introduces different time steps and hybrid state spaces consisting of logical components and continuous components.
+This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to safety constraints.
+A state-dependent fading channel is modeled to capture interference among transmission links, as well as the effects of mobile agents’ movements on successful wireless transmission.
+In order to address the heterogeneous system dynamics, the optimal control problem is formulated as the optimal constrained set stabilization of the MAS by establishing a necessary and sufficient condition for the Lyapunov-like performance of WCS with the expected decay rates.
+Using the semi-tensor product of matrices, a constrained optimal state transition graph is constructed to encode the constrained system dynamics as well as objective function, which further reduces the problem into a minimum-mean cycle problem for the graph.
+By studying the properties of the graph, the feasibility is proven, and an effective algorithm is proposed for the construction of optimal input sequences.
 An illustrative example is provided to demonstrate effectiveness of the proposed method.
 keywords:
-Constrained set stabilization; heterogeneous Industrial Internet of Things system; infinite-horizon optimal control; shared wireless medium.
+Heterogeneous system; infinite-horizon optimal control; shared wireless medium; constrained set stabilization; semi-tensor product of matrices.
 thanks: The work was supported by National Key R&D Program of China under the grant 2022YFB3303900 and the NSF of China under the grants 61933009, 62103272.
@@
 1.1 Background and motivation
-Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields including smart manufacturing through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011].
+Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields, including smart manufacturing, through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011].
 These plants consist of multiple physical systems with sensing, computation, communication and actuation capabilities, which flexibly interact with one another and human operators [Ahlen2019, vitturi2019].
 Due to such complex interactions among the entities, wired communication will reach its limits.
-Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming wireless control systems (WCSs).
-This abundance of wireless devices spawns the need of efficiently sharing the available wireless medium [gatsis2018].
-An integral aspect of IIoT in smart manufacturing is the use of mobile agents [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, forming a heterogeneous IIoT system, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019].
-In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as collision and dangerous position avoidance.
-Moreover, positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over, leading to variations in the measured channel gains [Agrawal2014, Quevedo2013].
-In order to ensure the performance of WCSs, it is necessary to explicitly examine the influence of mobile agents’ positions on channel states of the WCS, and design constrained agents motion strategy such that the desired performance requirements is guaranteed for the WCS.
-Such coupling brings heterogeneity to IIoT systems, namely, the WCS is coupled with the mobile agent system (MAS), which is a distinctive feature of IIoT systems compared with traditional WCSs.
-Some results considered the dependence of channel state on physical states, which mainly focused on channel characterization, and communication and controller design of WCSs [Agrawal2014, Ahlen2019, Kashiwagi2010].
-Few works addressed the heterogeneity feature induced by this dependence.
-A state-dependent channel model was proposed in [hu2019] to characterize the WCS and MAS coupling, and approximate optimal control and transmission power policies were obtained, under which the WCS achieves both safety and efficiency.
-This work was generalized in [hu2022] by considering a state-dependent Markov channel model to address time varying data rates, where optimal strategies were obtained by solving a constrained polynomial optimization problem.
-In these results, the state and input constraints for the MAS were ignored.
-In addition, only a simplified model consisting of a WCS with a single control loop and a single agent was considered.
-How to solve the coordination problem of WCSs and MASs in IIoTs with multiple plants being controlled over a shared wireless channel is largely unexplored.
-The main challenge is to model the fading channel which captures both the interference among transmission links, and the coupling between WCS and MAS, based on which to address the heterogeneity of the whole IIoT system.
+Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming multiloop wireless control systems (WCSs).
+An integral aspect of IIoT in smart manufacturing is the use of mobile agents, e.g., automated guided vehicles (AGVs) [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019].
+In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as moving within task areas and velocity limits while avoiding dangerous positions [Baumann2021].
+Moreover, the positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over [Agrawal2014, Quevedo2013].
+On-site measurements in factories reveal that the channel fading caused by machinery movements exhibits a piecewise behavior, i.e., the fading channel gain remains stationary as the machinery moves within a certain area [Ahlen2019].
+And each area is associated with a certain level of shadowing effects.
+This corresponds to a high level dynamics of MAS exhibiting logical characteristics.
+Additionally, the movements of mobile agents among areas after spatial discretization is typically on the time scale of seconds, while the wireless channel transmits packets on the order of milliseconds.
+This means that the coordination of the WCS and mobile agent system (MAS) introduces different time steps and hybrid state spaces, consisting of the continuous component of industrial plants and the logical component of mobile agents.
+These distinctive features make the modeling and analysis tools on traditional WCSs fail in addressing the IIoT system considered in the paper.
+Due to the complex interactions between the WCS and MAS, in order to ensure the control performance of the WCS, it is necessary to explicitly examine the influence of mobile agents’ positions on the channel states of the WCS, and to design a constrained motion strategy for agents such that the desired performance requirements are guaranteed for the WCS.
 1.2 Related work
@@
 For a survey on design and optimization for WCSs, please refer to [Park2018].
-As a part of IIoT, WCSs with fading channels have been extensively studied in the past few decades.
+As important components of IIoT systems, the MAS and WCS are typically considered separately in the literature.
+Recently, the finite-field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014].
+The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011].
+It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints.
+Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems.
+Noting that all the above results are established for finite-valued systems, i.e., the dynamics of all nodes in the system are the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to control performance requirements constraints.
+As for the other part of IIoT, WCSs with fading channels have been extensively studied in the past few decades.
 It is noted that the channel gains used to characterize shadow fading are traditionally modeled as independent identical distributed random processes [Gatsis2014] or Markov chains [zhangq1999], where the network state is assumed to be independent from physical states.
@@
 References [gatsis2015] and [gatsis2018] considered shared fading channels, where centralized and distributed channel-aware access mechanisms were respectively presented to meet the Lyapunov-like performance requirements while minimizing the total expected power consumption.
-It is obvious that these methods on performance analysis of WCS with fading channels fail in the IIoT scenario where the WCS and MAS coupling exists, not to mention the case of constrained MAS.
+The above channel models are clearly inadequate for characterizing the complex interactions between the WCS and MAS.
+In addition, all these results focus on the control performance of plants with only continuous state space.
+It is obvious that these methods fail in analyzing IIoT systems with hybrid state spaces considered in this paper.
 Few works addressing dependence of network state on physical states have been conducted in IIoT.
-A state-dependent bursty fading channel model was proposed in [hu2013], based on which distributed switching control [hu2015] and power control schemes [hu2020] were respectively proposed to ensure the performance of vehicular WCSs.
+A state-dependent bursty fading channel model was proposed in [hu2014].
 However, only homogeneous systems without constraints were considered in these works.
@@
 By solving a constrained polynomial optimization problem, optimal design strategies were obtained.
-In the above works, safety constraints of the mobile agents were ignored.
-In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, and coordinate with multiple mobile agents to complete manufacturing tasks, the proposed state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
-As for the other part of IIoT, recently, the finite field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014].
-The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011].
-It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints.
-Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems.
-The optimal problem with a set stabilization constraint was considered in [gao2021], where only the finite-horizon optimization until reaching the target set was addressed.
-Noting that all the above results are established for homogeneous finite-valued systems, i.e., the dynamics of all nodes in the system is the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to performance requirements constraints.
+In the above works, a common time step is adopted for the evolution of the MAS and WCS.
+Noting that the wireless channel typically transmit packets on the order of milliseconds, this requires mobile agents to move on a millisecond time scale, which is often unrealistic.
+In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, the state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
 1.3 Contributions
-In this paper, we focus on the optimal control of MAS with state and input constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling.
-A preliminary version of the paper appears in [wangs2022], where only controller design problem is addressed for the MAS to ensure the performance of WCS.
+In this paper, we focus on the optimal control of MAS with safety constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling.
+A preliminary version of the paper appears in [wangs2022], where a common time step is considered for the WCS and the leader-follower MAS.
 The main contributions of the paper are summarized as follows.
-First, to take into account the level of shadowing effects on channel gains of the WCS, the finite field network architecture is used to model the operational control for the MAS.
-The state-dependent fading channel is then modeled to characterize the dependence of wireless transmission of the WCS on mobile agents’ positions, leading to the state-dependent Lyapunov-like performance requirement of the WCS.
+  • We propose a heterogeneous IIoT system framework to capture the coupling between the WCS and MAS, where the multiloop WCS is described as a switched dynamics model depending on wireless transmission, and the finite-field architecture is used to model the high level dynamics of MAS moving among areas with different levels of shadow effects.
+Different time steps are adopted to capture different time scales between the wireless transmission of control systems and the movements of mobile agents among discretized areas.
+Effect of mobile agents’ movements on successful wireless transmission is characterized by a state-dependent fading channel, which captures the spectrum resource competition and interference among different loops.
 Based on this model, the infinite-horizon optimal control problem is established for the heterogeneous IIoT system to quantify the tradeoff between WCS’s power consumption and MAS’s control cost, while ensuring the Lyapunov-like performance requirement of the WCS and the safety constraints of the MAS.
-Second, optimal input sequences are designed for the MAS to solve the infinite-horizon optimal control problem.
-A criterion in terms of constrained set stabilization for the MAS is established to ensure the Lyapunov-like performance of WCS, which equivalently converts the optimal control problem into the infinite-horizon optimal constrained set stabilization problem of the MAS.
-Since each agent updates its states from a finite field, it is difficult to apply classical nonlinear system theory to solve the optimal problem.
-We adopt the semi-tensor product approach to represent the MAS as a discrete-time bilinear system, based on which a necessary and sufficient condition is presented for the constrained set stabilization of MAS.
-In order to address state and input constraints of the MAS, a constrained optimal state transition graph is constructed, by analyzing which the feasibility of the considered optimal control problem is proved.
-Then, an algorithm is proposed to design optimal input sequences by resorting to minimum-mean cycles in weighted graph.
+  • To address the hybrid nature of the heterogeneous IIoT system, we first equivalently convert the optimal control problem into the optimal constrained set stabilization problem of the MAS.
+In order to address the nonlinearity due to the finite-field model of the MAS, we adopt the semi-tensor product approach, and construct a constrained optimal state transition graph to handle safety constraints elegantly.
+By encoding the constrained system dynamics and objective function into the graph, the problem is further deduced into a minimum-mean cycle problem on the constrained optimal state transition graph.
+Then, through studying the properties of the graph, we propose an efficient algorithm to design optimal input sequences that ensure wireless control performance and safety constraints.
 This paper is organized as follows.
 In Section 2, we characterize the heterogeneous IIoT system and formulate the problem studied in this paper.
-Section 3 considers Lyapunov-like performance analysis of the WCS.
-Feasibility analysis and optimal controller design are proceeded in Section 4 and Section 5, respectively.
+Section 3 formulates the optimal control problem as the optimal constrained set stabilization problem for the MAS.
+Design of optimal controllers using the graphical method is provided in Section 4.
 Section 5 demonstrates main results by an illustrative example, which is followed by the conclusion in Section 6.
 The key notations used in this paper are summarized in Table 1.
-Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$ [chengqi2011, cheng2011].
+Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$, and $\otimes$ is the Kronecker product of matrices [chengqi2011, cheng2011].
 Since the semi-tensor product is a generalization of the conventional matrix product, the symbol $\ltimes$” is omitted in most places of this paper when no confusion arises.
@@
 Notations Definitions
-$\mathbb{N}_{+}$ Set of natural numbers
 $\mathcal{D}_{s}$ Logic domain $\{0,1,\cdots,s-1\}$
-$\mathcal{D}_{n}^{m}$ Cartesian product $\underbrace{\mathcal{D}_{n}\times\cdots\times\mathcal{D}_{n}}_{m}$
 $I_{s}$ $s$-dimensional identity matrix
@@
 $[F]_{i,j}$ $(i,j)$-th entry of matrix $F$
-$\delta_{s}[i_{1}\;i_{2}\;\cdots\;i_{t}]$ Logical matrix $F$ with $[F]_{:,j}=\delta_{s}^{i_{j}}$
-${\mathcal{L}}^{n\times m}$ Set of $n\times m$ logical matrices
 $\mathbb{R}^{n\times m}$ Set of $n\times m$ real matrices
@@
 $\ltimes$ Semi-tensor product
-$\otimes$ Kronecker product
+[…]
+${}\mathbb{P}\{\cdot\}$ Probability of a random event
+$\mathbb{E}\{\cdot\}$ Expectation of a random variable
 $|\mathcal{M}|$ Cardinal number of set $\mathcal{M}$
-$\mathcal{M}\setminus\mathcal{N}$ Set $\{x\in\mathcal{M}:x\notin\mathcal{N}\}$
 2 System model and problem formulation
-Figure 1: Framework of a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a factory floor, discretized into $\kappa$ subregions.
-‘S’ and ‘A’ respectively represent sensing and actuation capabilities.
-The local channel state of each control loop $i$ ($\gamma_{i}\in\mathcal{D}_{r}$) is dependent on positions of agents $\alpha_{j}\in\mathcal{D}_{\kappa}$, $j=1,\cdots,n$.
 We consider a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a typical smart manufacturing scenario (Fig.
@@
 In this process, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over, resulting in the coupling between the WCS and MAS.
-According to the level of shadowing effects on channel gains of the WCS, partition all possible positions of the mobile agents in the factory floor into $\kappa$ overlapping sets.
-Each set corresponds to a certain level of shadowing effects.
-Correspondingly, the whole workspace is discretized into $\kappa$ overlapping planar subregions, called cells $\{0,1,\cdots,\kappa-1\}$.
+Figure 1: Framework of a heterogeneous IIoT consisting of a multiloop WCS and an MAS in a factory floor.
+‘S’ and ‘A’ respectively represent sensing and actuation capabilities.
+The workspace is partitioned into $\kappa$ cells $\{0,1,\cdots,\kappa-1\}$, and each cell is abstracted as a logical state indicating a level of shadowing effects.
+Each mobile agent $j$ moves within its task area, leading to logical state transitions among cells in the state constraint set $\mathcal{C}_{\alpha_{j}}$.
 2.1 Plants
-Denote the state of plant $i$ at time $k$ by $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
-$x_{i}(k+1)=A_{i}x_{i}(k)+B_{i}v_{i}(k)+\xi_{i}(k),$ (1)
-where $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}\times 1}$ is the driving input, and $\{\xi_{i}(k):k\in\mathbb{N}\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
+Denote the state of plant $i$ at time instant $t_{l}$ by $x_{i}(l)\in\mathbb{R}^{n_{i}}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
+$x_{i}(l+1)=A_{i}x_{i}(l)+B_{i}v_{i}(l)+\xi_{i}(l),$ (1)
+where $\{t_{l}\}_{l=0}^{\infty}$ with $t_{l}<t_{l+1}$ is the sequence of sampling time instants, $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}}$ is the driving input, and $\{\xi_{i}(l):l=0,1,\cdots\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
 2.2 Sensors and access point
 $q$ sensors are placed to monitor the plants.
-At every slot $k$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not adapting to its local wireless channel state.
-We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at slot $k$ is $\gamma_{i}(k)\in\mathcal{D}_{r}$.
-Assume that the wireless communication policy is pre-designed as
-$\varphi_{i}(k)=h_{i}(\gamma_{i}(k)),$ (2)
-where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(k)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at slot $k$.
-In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot.
+At every time instant $t_{l}$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not.
+We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at time instant $t_{l}$ is $\gamma_{i}(l)\in\mathcal{D}_{r}$.
+Assume that the wireless communication policy111 Regardless of the effect of mobile agents’ movements on wireless transmission, following the approach in [gatsis2018], it can be proven that the optimal transmission policies ensuring the Lyapunov-like control performance of the WCS only depend on the channel states $\gamma_{i}(l)$.
+is pre-designed as
+$\varphi_{i}(l)=h_{i}(\gamma_{i}(l)),$ (2)
+where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(l)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at time instant $t_{l}$, and each sensor transmits only when its corresponding local channel state is favorable enough and backs off otherwise.
+In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the local channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot.
 The pilot signal may also serve as a synchronization purpose [gatsis2018, hu2011].
@@
 Then, control decisions are wirelessly fed back to actuator $i$.
-These $q$ independent control loops form a multiloop WCS.
+These $q$ control loops form a multiloop WCS.
 2.3 Multiloop WCS
 The transmission of measurements and control decisions might fail due to packet decoding errors and packet collisions.
-In this paper, we focus on the uplink transmission from the sensors to the access point, the idea can be extended into the downlink transmission scenario straightforwardly.
-Use $\lambda_{i}(k)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time $k$.
-Then, the measurement $x_{i}(k)$ is used to compute the control input $v_{i}(k)$ by pre-designed control policy at a successful transmission ($\lambda_{i}(k)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$.
-Otherwise, an open-loop system with system matrix $A_{o,i}$ which may be unstable is obtained.
-The resulting system can be uniformly described as the following switched model:
-$x_{i}(k+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=1,\\ A_{o,i}x_{i}(k)+\xi_{i}(k),&\lambda_{i}(k)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
+Assume that both wireless uplink and downlink have the same channel characteristics [hu2019].
+In this case, we focus on the uplink transmission from the sensors to the access point, and the idea can be extended into the downlink transmission scenario straightforwardly.
+Use $\lambda_{i}(l)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time instant $t_{l}$.
+Then, the measurement $x_{i}(l)$ is used to compute the control input $v_{i}(l)$ at a successful transmission ($\lambda_{i}(l)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$.
+Otherwise, an open-loop system with the system matrix $A_{o,i}$, which may be unstable, is obtained.
+The resulting system can be described as the following switched model:
+$x_{i}(l+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=1,\\ A_{o,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
 2.4 Finite-field MAS
-Due to the discretization of the workspace, it is the cell that the agents are located in, not the specific position, that directly affects the channel condition of the WCS.
-In this way, the state of an agent is defined as the index of the cell that the agent is located in, i.e., the state space of each agent is the finite set $\mathcal{D}_{\kappa}$.
-At each time step, each agent updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite field model:
-$\alpha_{j}(k+1)=\{_{\kappa}\sum\limits_{l\in\bar{\mathcal{I}}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
-where ¯$\bar{\mathcal{I}}_{j}=\mathcal{I}_{j}\cup\{j\}$, $\mathcal{I}_{j}$ is the set of in-neighbors of $j$, $\alpha_{j}(k)$, $u_{j}(k)$ respectively denote the state and input of $j$ at time $k$, $a_{j,l}\in\mathcal{D}_{\kappa}$ are the weights, and operations $+_{\kappa}$, ×$\times_{\kappa}$ denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$, respectively [Lidl1996].
-In this way, the state $\alpha_{j}(k)$ will evolve over the finite cells $\{0,1,\cdots,\kappa-1\}$.
-Such finite field architecture has been widely studied to address problems for networks of agents with finite state space [Pasqualetti2014, Sundaram2013].
-To ensure production safety such as collision and dangerous position avoidance, we impose state constraints $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}^{n}$ and state-dependent input constraints $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))\subseteq\mathcal{D}_{\kappa}^{n}$.
-This is one of the differences with the existing results [hu2022, hu2019] which model the mobile agent by an unconstrained Markov decision process.
-The existence of constraints makes the control of MAS much more complex.
+An MAS coordinates with the multiloop WCS to jointly perform smart manufacturing tasks.
+Mobile agents are driven by the standard single-integrator dynamics222 This paper uses the model as an illustration of how to abstract the finite-field MAS from the physics dynamics.
+The proposed results are general and can be extended to other models for mobile agents.
+as
+$z_{j}(k+1)=z_{j}(k)+T\nu_{j}(k),j=1,\cdots,n,$
+where $z_{j}(k),\nu_{j}(k)\in\mathbb{R}^{2}$ respectively denote the position and velocity of agent $j$ in the two-dimensional space at time instant $t_{\tau k}$, $\{t_{\tau k}\}_{k=0}^{\infty}$ is the sequence of sampling time instants with time step $T$, and $\tau$ is a positive integer indicating different time steps for the evolution of the MAS and WCS.
+When coordinating with the WCS to jointly perform overall tasks, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over.
+Note that what matters for shadow fading is the position in space within certain intervals [Ahlen2019], not the specific position determined by the single-integrator dynamics.
+Partition the whole workspace into $\kappa$ non-overlapping cells $\{0,1,\cdots,\kappa-1\}$.
+Abstract the cell in which each mobile agent is located as a logical state in $\mathcal{D}_{\kappa}$.
+Each agent interacts with others by on-board sensors, and updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite-field model:
+$\alpha_{j}(k+1)\!=\!\{_{\kappa}\!\sum\limits_{l\in\mathcal{I}_{j}\cup\{j\}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
+where $\mathcal{I}_{j}$ is the set of all in-neighbors of $j$, $a_{j,l}\in\mathcal{D}_{\kappa}$ denotes the weight, operations $+_{\kappa}$ and ×$\times_{\kappa}$ respectively denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$ [Lidl1996], $\alpha_{j}(k)$, $u_{j}(k)\in\mathcal{D}_{\kappa}$ respectively denote the state and control input of agent $j$ at time instant $t_{\tau k}$.
+The feedback term $\{_{\kappa}\sum_{l\in\mathcal{I}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)$ together with $u_{j}(k)$ can describe several mobile agents’ controllers, including consensus, rendezvous and formation [Mesbahi, Pasqualetti2014, Sundaram2013].
+The constraint on state $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}$ with $\alpha_{j}(k)\in\mathcal{C}_{\alpha_{j}}$, $j=1,\cdots,n$ is imposed to ensure each agent $j$ moves within its task areas $\mathcal{C}_{\alpha_{j}}$ while avoiding dangerous positions, such as collisions with industrial plants.
+In addition, the transition constraint $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))$ is imposed to restrict agents’ movements within the actual velocity limits.
+Such safety constraints distinguish the model from previous works [hu2022, hu2019] representing the mobile agent by an unconstrained Markov decision process.
 2.5 State-dependent fading channel model
-In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and thus influence the local channel state.
-We call a packet collision happens on system $i$, rendering system $i$’s packet decoding failed, if some other sensor $j\neq i$ transmits in the same time slot.
-If sensor $i$ transmits and has a collision-free time slot, the success of decoding the packet at the access point depends on the local channel state of link $i$, then the state of the MAS.
-Such WCS and MAS coupling is modeled by the following state-dependent fading channel:
-$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\lambda}_{i}(\alpha),$ (5)
-where $\bar{\lambda}_{i}(\alpha):\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$, $\lambda_{i}(k)=1$ indicates the successful transmission of system $i$.
+In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and finally influence the probability of successful wireless transmission.
+With different time steps, during each sampling interval $[t_{\tau k},t_{\tau(k+1)})$ of the MAS, there are $\tau$ time instants of the wireless transmission as $t_{l}$, $l={\tau k},{\tau k+1},\cdots,{\tau(k+1)-1}$, where the $k$ is determined by the relation $k=\lfloor l/\tau\rfloor$.
+This means that the successful transmission probabilities for WCS during these $\tau$ consecutive time instants $t_{l}$ depend on the state of MAS at time instant $t_{\tau k}$, i.e., $\alpha(k)$.
+Then, the WCS and MAS coupling is modeled by the following state-dependent fading channel:
+$\mathbb{P}\Big{\{}\lambda_{i}(l)=1|\alpha(k)=\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}=\bar{\lambda}_{i}(\alpha),$ (5)
+where $\lambda_{i}(l)=1$ indicates the successful transmission of link $i$ at time instant $t_{l}$.
+Let $\eta_{i}(l)=1$ denote the successful packet decoding of link $i$. As shown in Fig. 2, with wireless communication policy (2), it holds
+$\bar{\lambda}_{i}(\alpha)\!=\!\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)\mathbb{P}\{\eta_{i}(l)\!=\!1|\alpha(k)\!=\!\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\},$
+where $\bar{\gamma}_{i}(c,\alpha)=\mathbb{P}\{\gamma_{i}(l)=c|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$ characterizes the probability distribution for the local channel state, and can be easily obtained by measuring the local fading amplitude for each MAS state by the statistical methods in [Agrawal2014, Kashiwagi2010].
+Taking into account the interference among transmission links by adopting the collision channel model [gatsis2018], the last term in $\bar{\lambda}_{i}(\alpha)$ is equal to $\mathbb{P}\{\hbox{SNR}_{i}(l)\geq\eta_{i}^{0}|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor,\varphi_{i}(l)=1,\varphi_{j}(j)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(l)=0|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$, where $\hbox{SNR}_{i}(l)$ and $\eta_{i}^{0}$ denote the signal-to-noise ratio and the specified decodability threshold, respectively.
+Figure 2: State-dependent fading channel (5).
 Eq.
 (5) presents a general state-dependent channel model.
-Communication policies for sensors and the interference among transmission links can be embedded into this model.
 In this model, the probability of a successful transmission for WCS (3) depends on the state of MAS (4).
-It is quite different from the traditional shadow fading model [Gatsis2014, zhangq1999] assuming the channel states independent from the physical states.
-Such coupling makes the traditional methods on performance analysis of WCSs [gatsis2015, gatsis2018, Hristu2001, le2011, Molin, zhang2006] fail in the scenario in this paper.
+In addition, as shown in Fig.
+2, Eq.
+(5) takes into account wireless communication policies, the interference among transmission links, and different time steps for the WCS and MAS, which cannot be captured by the existing channel models in [hu2022, hu2019].
 2.6 Problem formulation
-Considering the state-dependent fading channel model (5), given initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$ of the MAS, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ for MAS (4) such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k\in\mathbb{N}$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
+Considering the state-dependent fading channel model (5), given an initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ for MAS (4) such that $\alpha(k)\in\mathcal{C}_{\alpha}$, $k=0,1,\cdots$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
 Definition 1.
-The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}\subseteq\mathcal{D}_{\kappa}^{n}$ and a positive integer $K$ such that the inequalities
-$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]\leq\rho_{i}V_{i}(x_{i}(k))$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+Tr(Q_{i}\Xi_{i}),\forall\ x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ (6)
-hold for any $k\geq K$, any $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$ and any $i\in\{1,\cdots,q\}$, where $\mathbb{E}$ is the expectation, $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
-In (6), $\rho_{i}$ represents desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation.
+The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}$ and a positive integer $L_{0}$ such that the inequalities
+$\mathbb{E}\Big{[}V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{]}$
+$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\rho_{i}V_{i}(x_{i}(l))+Tr(Q_{i}\Xi_{i}),\forall x_{i}(l)\in\mathbb{R}^{n_{i}}$ (6)
+hold for any $l\geq L_{0}$, any $x_{i}(l)\in\mathbb{R}^{n_{i}}$ and any $i\in\{1,\cdots,q\}$, where $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
+In (6), $\rho_{i}$ represents the desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation.
 Similar kinds of performance requirements as in (6) have been adopted for WCSs [gatsis2015, gatsis2018].
-However, different from them, Definition 1 explicitly explores the state-dependent fading channel.
+Different from them, Definition 1 explicitly explores the effect of mobile agents’ movements on the successful wireless transmission of the WCS.
+In addition, it allows adopting different time steps to capture different time scales between the wireless transmission of control systems and the movements of mobile agents from cell to cell.
 Our second goal in this paper is to determine optimal input sequence ${\bf{u}}^{\ast}$ such that the average cost function
-$J_{{\bf{u}}}(\alpha_{0})=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\sum\limits_{k=0}^{K-1}\Big{\{}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda g(\alpha(k;\alpha_{0},{\bf{u}}),u(k))\Big{\}}$ (7)
-reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}^{m+n}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
-$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
+$J_{{\bf{u}}}=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\Big{\{}\sum\limits_{l=0}^{K-1}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\Big{\{}\varphi_{i}(l)=1|\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}$
+$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda\sum\limits_{k=0}^{\lfloor(K-1)/\tau\rfloor}g(\alpha(k),u(k))\Big{\}}$ (7)
+reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\mu_{i}$ is the transmit power for sensor $i$, $\sum_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS, and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
+$P_{1}:$ $\min_{{\bf{u}}}J_{{\bf{u}}}$
 […]
-$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},$
-$~{}~{}~{}~{}~{}~{}~{}~{}u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (8)
-In the above problem, there are heterogeneous system dynamics constraints (3), (4), WCS’s performance requirements (6) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, which make the resolution of this problem difficult.
-Figure 2: Main idea of solving the infinite-horizon optimal control problem (8).
-3 Lyapunov-like performance analysis over the state-dependent fading channel
-In the optimal control problem (8), Lyapunov-like performance requirements constraints (6) brings about couplings between the WCS and the MAS, making the resolution of problem (8) challenging.
-To address this issue, we convert constraint (6) for WCS (3) and safety constraints to the constrained set stabilization of MAS (4) using the state-dependent fading channel (5).
-In this way, all constraints in problem (8) are related to the MAS, which facilitates its optimal input sequence design.
-Motivated by this observation, the main idea of solving problem (8) is illustrated in Fig.
-2.
-3.1 Characterization of state-dependent fading channel
-In this part, we characterize the state-dependent fading channel in (5) by taking into account of the wireless communication policy and the interference among transmission links.
-For fixed wireless communication policy, the probability of successful transmission for link $i$ can be expressed as $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k)=\alpha\}=\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}$, where $\eta_{i}(k)=1$ indicates the successful packet decoding.
-Due to the discrete channel state, the local channel state in the presence of shadow fading caused by movements of mobile agents is computed as follows:
-$\mathbb{P}\{\gamma_{i}(k)=a|\alpha(k)=\alpha\}=\bar{\gamma}_{i}(a,\alpha),i=1,\cdots,q,$ (9)
-where $\mathbb{P}$ is the probability, and $\bar{\gamma}_{i}:\mathcal{D}_{r}\times\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$ can be easily obtained by measuring the local fading amplitude for each MAS state according to the statistical methods in [Agrawal2014, Kashiwagi2010] and reference therein.
-Then, according to (2), we have $\mathbb{P}\{\varphi_{i}(k)=1|\alpha(k)=\alpha\}=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)$
-Since the transmit power is fixed, the probability that the packet of system $i$ is successfully decoded at the access point depends on the local channel state of link $i$, then the state of the MAS as
-$\mathbb{P}\{\eta_{i}(k)=1|\alpha(k)=\alpha\}=\bar{\eta}_{i}(\alpha),$ (10)
-where $\bar{\eta}_{i}:\mathcal{D}_{\kappa}^{n}\rightarrow[0,1]$. The interference among transmission links is taken into account by adopting the collision channel model. Then, it holds
-$\bar{\eta}_{i}(\alpha)=$$\mathbb{P}\{\hbox{SNR}_{i}(k)\geq\eta_{i}^{0}|\alpha(k)=\alpha,\varphi_{i}(k)=1,$
-$\varphi_{j}(k)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(k)=0|\alpha(k)=\alpha\},$
-where $\hbox{SNR}_{i}(k)=\mu_{i}\gamma_{i}^{2}(k)/N_{0}$, $N_{0}$ and $\eta_{i}^{0}$ respectively denote the signal-to-noise ratio [hu2019], noise power and the specified decodability threshold of system $i$.
-Based on the above analysis, we obtain the following result.
-Lemma 2.
-For the state-dependent fading channel (5), it holds $\bar{\lambda}_{i}(\alpha)=\sum_{a=0}^{r-1}\bar{\gamma}_{i}(a,\alpha)h_{i}(a)\bar{\eta}_{i}(\alpha)$, $\alpha\in\mathcal{C}_{\alpha}$, $i=1,\cdots,q$.
-3.2 Lyapunov-like performance analysis of WCS
-Based on the characterization of state-dependent fading channel, we can now analyze the Lyapunov-like performance of the WCS.
-For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})):k\in\mathbb{N}\}$ an admissible input sequence in the following.
+$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(0)=\alpha_{0},$
+$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k)\in\mathcal{C}_{\alpha},u(k)\in\mathcal{C}_{u}(\alpha(k)),k=0,1,\cdots.$ (8)
+3 Optimal wireless control as optimal constrained set stabilization
+In the optimal control problem $P_{1}$, there are heterogeneous system dynamics constraints consisting of continuous component (3) and logical component (4) with different time steps.
+In addition, control performance (6) of the continuous system is constrained by the logical one.
+The disparity in analyzing these distinct systems makes the resolution of the problem difficult.
+To address these challenges, we convert WCS’s performance requirements (6) together with safety constraints (8) into the constrained set stabilization of MAS (4) using the state-dependent fading channel (5).
+In this way, problem $P_{1}$ is converted to the optimal constrained set stabilization problem of the MAS.
+For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ an admissible input sequence in the following.
 Definition 2.
-Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{M}$ holds for any $k\geq K_{0}$.
+Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k)\in\mathcal{M}$ holds for any $k\geq K_{0}$.
 Definition 3.
-A subset $\mathcal{S}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset (CCIS) of MAS (4), if for any $\alpha(0)=\alpha_{0}\in\mathcal{S}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{S}$ holds for any $k\in\mathbb{N}$.
-Obviously, the union of any two CCISs of MAS (4) is another CCIS.
-The union of all CCISs contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
+A subset $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset of MAS (4), if for any $\alpha_{0}\in\mathcal{M}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k)\in\mathcal{M}$ holds for any $k=0,1,\cdots$.
+Obviously, the union of any two constrained control invariant subsets of MAS (4) is another constrained control invariant subset.
+The union of all constrained control invariant subsets contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
 Theorem 4.
-The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints $\mathcal{C}_{\alpha}$, $\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}}))$, if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
-$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}\times 1},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (11)
+The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints (8), if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
+$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (9)
 Proof 3.1.
-For each subsystem $i$, since $\mathbb{E}[\xi_{i}(k)]=\mathbf{0}$ and $\lambda_{i}(k)$ is independent of $x_{i}(k)$, by (3), it holds
-$\mathbb{E}[V_{i}(x_{i}(k+1))|x_{i}(k),\alpha(k;\alpha_{0},{\bf{u}})]$
-$=$ $\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(k)$
-+$+\mathbb{P}\{\lambda_{i}(k)=0|\alpha(k;\alpha_{0},{\bf{u}})\}x_{i}^{\top}(k)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(k)$
+For each subsystem $i$, since $\mathbb{E}[\xi_{i}(l)]=\mathbf{0}$ and $\lambda_{i}(l)$ is independent of $x_{i}(l)$, by (3), it holds
+$\mathbb{E}[V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\lfloor l/\tau\rfloor]$
+$=$ $\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(l)$
++$+\mathbb{P}\{\lambda_{i}(l)=0|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(l)$
 $+T_{r}(Q_{i}\Xi_{i}).$
-Then, condition (6) subject to the multiloop WCS dynamics (3) is equivalent to the condition that for all $x_{i}(k)\neq\mathbf{0}$, it holds
-$\mathbb{P}\{\lambda_{i}(k)=1|\alpha(k;\alpha_{0},{\bf{u}})\}$
-$\geq$ $\frac{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(k)}{x_{i}^{\top}(k)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(k)},$ (12)
-which is equivalent to $\bar{\lambda}_{i}(\alpha(k;\alpha_{0},{\bf{u}}))\geq s_{i}$ with $s_{i}$ given in (11), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
+Then, condition (6) subject to multiloop WCS dynamics (3) and safety constraints holds, if and only if there exist an admissible input sequence and a positive integer $L_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $x_{i}(l)\neq\mathbf{0}$, it holds
+$\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$
+$\geq$ $\frac{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(l)}{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(l)},$ (10)
+which are equivalent to the condition that there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$, such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $k\geq K_{0}$, it holds $\bar{\lambda}_{i}(\alpha(k))\geq s_{i}$, $s_{i}$ is given in (9), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
 Next, we prove that MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$ if and only if it is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$.
@@
 This is proved by a contradiction argument.
-Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence $\bar{{\bf{u}}}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},\bar{{\bf{u}}})\in\Omega(s)$.
-Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}})\notin I(\Omega(s))$.
-Since $I(\Omega(s))$ is a largest CCIS, for any $\bar{\alpha}\in\Omega(s)\setminus I(\Omega(s))$ and any admissible input sequence ${\bf{u}}$, there exists an integer $k(\bar{\alpha},{\bf{u}})$ such that $\alpha(k(\bar{\alpha},{\bf{u}});\bar{\alpha},{\bf{u}})\notin\Omega(s)$.
-Then, it holds
-$\alpha(k_{0}+k(\alpha(k_{0};\alpha_{0},\bar{{\bf{u}}}),\tilde{{\bf{u}}});\alpha_{0},\bar{{\bf{u}}})\notin\Omega(s),~{}\tilde{u}(k)=\bar{u}(k+k_{0}),$
-which is a contradiction. Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. $\Box$
-By (12), $s_{i}$ represents the lower bound of the probability of successful transmission for system $i$ that ensures the desired Lyapunov-like performance given in Definition 1 under the WCS and MAS coupling.
+Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence ${\bf{u}}^{\prime}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\Omega(s)$.
+Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then under input sequence ${\bf{u}}^{\prime}$, there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0})\notin I(\Omega(s))$.
+Since $I(\Omega(s))$ is a largest constrained control invariant subset, for any $\alpha^{\prime}\in\Omega(s)$ and any admissible input sequence ${\bf{u}}$, if $\alpha^{\prime}\notin I(\Omega(s))$, there exists an integer $k(\alpha^{\prime},{\bf{u}})$ such that $\alpha(k(\alpha^{\prime},{\bf{u}}))\notin\Omega(s)$ from $\alpha^{\prime}$.
+Then, under input sequence ${\bf{u}}^{\prime}$, it holds $\alpha(k_{0}+k(\alpha(k_{0}),{\bf{u}}^{\prime\prime}))\notin\Omega(s)$ from $\alpha_{0}$, $u^{\prime\prime}(k)=u^{\prime}(k+k_{0})$, which is a contradiction.
+Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$.
+$\Box$
+By (10), $s_{i}$ is the lower bound of the successful transmission probability for system $i$ ensuring the desired Lyapunov-like performance in Definition 1.
+The value of $s_{i}$ can be obtained by the following linear program:
+$\min_{\theta_{i}\geq 0}~{}\theta_{i}$
+[…]
+which can be solved using the toolbox such as YALMIP in MATLAB.
 In addition, $\Omega(s)$ consists of all constrained states of the MAS satisfying the performance requirements constraint (6) of WCS (3).
-Thus, for the optimal control problem (8), all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable, that is, problem (8) is equivalently converted into the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
-$\min_{{\bf{u}}}J_{{\bf{u}}}(\alpha_{0})$
+For the optimal control problem $P_{1}$, all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable.
+Since $\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}=\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)$, ${\bf{u}}^{\ast}$ is a solution to $P_{1}$, if and only if it is a solution to the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
+$P_{2}:$ $\min_{{\bf{u}}}\bar{J}_{{\bf{u}}}=\lim_{K\rightarrow\infty}\frac{1}{K}\sum_{k=0}^{K-1}\bar{g}(\alpha(k),u(k))$
 […]
-$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k;\alpha_{0},{\bf{u}})\in I(\Omega(s)),\forall\ k\geq K_{0},$
-$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k;\alpha_{0},{\bf{u}})\in\mathcal{C}_{\alpha},\forall\ k<K_{0},$
-$~{}~{}~{}~{}~{}~{}~{}\,u(k)\in\mathcal{C}_{u}(\alpha(k;\alpha_{0},{\bf{u}})),\forall\ k\in\mathbb{N}.$ (13)
-4 Feasibility analysis
-In this section, we first reformulate the MAS in a tractable form, and then present a criterion on the constrained set stabilization of MAS (4).
-Finally, the feasibility of optimal control problem (13) is analyzed based on the graphical method.
-4.1 Algebraic reformulation of the MAS
-Due to the highly nonlinear finite field model, it is difficult to apply classical nonlinear system theory to investigate (4).
+$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k)\in I(\Omega(s)),\forall k\geq K_{0},$
+$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k)\in\mathcal{C}_{\alpha},\forall k<K_{0},$
+where $\bar{g}(\alpha(k),u(k))$ is the joint stage cost defined as
+$\bar{g}(\alpha(k),\!u(k))\!=\!\tau\!\sum_{i=1}^{q}\!\mu_{i}\!\sum_{c=0}^{r-1}\!\bar{\gamma}_{i}(c,\alpha(k))h_{i}(c)\!+\!\lambda g(\alpha(k),\!u(k)).$
+Let the optimal objective values of problems $P_{1}$ and $P_{2}$ be $J^{\ast}=J_{{\bf{u}}^{\ast}}$ and $\bar{J}^{\ast}=\bar{J}_{{\bf{u}}^{\ast}}$, respectively.
+It is easy to obtain that $J^{\ast}=\bar{J}^{\ast}/\tau$.
+4 Infinite-horizon optimal controller design
+Note that both the state space and control space of the MAS dynamics (4) are finite.
+This property allows us to adopt the graphical method to address the optimization problem $P_{2}$.
+By reformulating the logical-state MAS in a tractable form via the semi-tensor product approach, the constrained optimal state transition graph is firstly constructed to encode the constraints and the objective function of $P_{2}$.
+Based on this, the feasibility of problem $P_{2}$ is deduced to the existence of cycles in the graph, and the design of optimal input sequences is modeled as a minimum-mean cycle problem.
+An algorithm is then proposed to construct optimal controllers.
+4.1 Feasibility analysis via a graphical method
+Due to the highly nonlinear finite-field model, it is difficult to apply classical nonlinear system theory to investigate (4).
 Recently, the semi-tensor product approach has emerged as an elegant tool to deal with such issues [chengqi2011, cheng2011].
-The basic idea is to embed the nonlinear finite filed dynamics in (4) into a higher dimensional space where its evolution is linear.
-In this part, by virtue of the semi-tensor product, we convert MAS (4) into a bilinear equivalent algebraic form, which facilitates the sequential studies.
-Set
-$\bar{a}_{j,l}=\left\{\begin{array}[]{ll}a_{j,l},&\hbox{if}\ l\in\bar{\mathcal{I}}_{j},\\ 0,&\hbox{otherwise},\end{array}\right.$
-and represent $i-1\in\mathcal{D}_{\kappa}$ by the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$.
-Then, $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k),\bar{a}_{j,l}\in Col(I_{\kappa})$, respectively.
-Letting $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)$, we have $\mathcal{C}_{\alpha},\mathcal{C}_{u}(\alpha(k))\subseteq Col(I_{N})$, where $N:=\kappa^{n}$.
-Then, based on similar calculation to [wangs2022], we have the following result.
-Lemma 6.
-MAS (4) has the following equivalent algebraic form:
+The basic idea is to embed the nonlinear finite-filed dynamics in (4) into a higher dimensional space where its evolution is linear.
+Specially, representing $i-1\in\mathcal{D}_{\kappa}$ as the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$, then $\alpha_{i}(k),u_{i}(k)\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k)\in Col(I_{\kappa})$, respectively.
+According to similar calculation to [wangs2022], we convert MAS (4) into the following bilinear equivalent algebraic form:
 $\alpha(k+1)=Fu(k)\alpha(k),$ (11)
-where $F\in\mathcal{L}^{N\times N^{2}}$ satisfies $[F]_{:,l}=\ltimes_{j=1}^{n}[F_{j}]_{:,l}$, $l=1,\cdots,N^{2}$, $F_{j}=S_{\kappa}^{n}\ltimes_{l=1}^{n}[I_{\kappa^{l-1}}\otimes(M_{\kappa}\bar{a}_{j,l})]W_{[\kappa,\kappa^{n}]}({\bf{1}}_{\kappa^{j-1}}^{\top}\otimes I_{\kappa}\otimes{\bf{1}}_{\kappa^{n-j}}^{\top})$, $j=1,\cdots,n$, $S_{\kappa}$ and $M_{\kappa}$ respectively represent the structural matrices for $+_{\kappa}$ and ×$\times_{\kappa}$ [Liy2019], ${\bf{1}}_{s}$ represents the $s\times 1$ vector with all entries being $1$, and $W_{[s,t]}$ represents the swap matrix defined as $W_{[s,t]}=[I_{t}\otimes\delta_{s}^{1}~{}\cdots~{}I_{t}\otimes\delta_{s}^{s}]$.
-4.2 Criterion of constrained set stabilization
-In this part, based on the constrained reachable set, we present a criterion for the constrained set stabilization of MAS (14).
-Denote the set of constrained states reachable from $\alpha_{0}$ by $\mathcal{R}(\alpha_{0})$. Then, we have the following result.
-Lemma 7.
-MAS (14) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$.
+where $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)\subseteq Col(I_{N})$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)\subseteq Col(I_{N})$, $N:=\kappa^{n}$, and $F$ is a logical matrix [chengqi2011, cheng2011] characterizing all state-transition information of the MAS.
+Then, the set of admissible inputs steering MAS (11) from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ in one step can be obtained as
+$U_{a,b}=\{\delta_{N}^{l}\in\mathcal{C}_{u}(\delta_{N}^{a}):[F\delta_{N}^{l}]_{b,a}=1\}.$ (12)
+In addition, the set of optimal inputs enabling the transition with the lowest cost is
+$\bar{U}_{a,b}={\arg\min}_{\delta_{N}^{l}\in U_{a,b}}\bar{g}(\delta_{N}^{a},\delta_{N}^{l}).$ (13)
+Figure 3: Constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$.
+To determine the feasibility of optimization problem $P_{2}$, we construct the constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$ to encode constraints and the objective function.
+Regard each state $\delta_{N}^{a}\in\mathcal{C}_{\alpha}$ as a vertex on a directed graph, and encode each constrained state transition from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ into edge $(\delta_{N}^{a},\delta_{N}^{b})$.
+Additionally, assign the weight $w(\delta_{N}^{a},\delta_{N}^{b})=\bar{g}(\delta_{N}^{a},\delta_{N}^{l})$ to the edge $(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}$ to encode the lowest joint cost under the optimal input $\delta_{N}^{l}\in\bar{U}_{a,b}$.
+See Fig.
+3 for an illustration.
+For an input sequence ${\bf{u}}$, we say that the state trajectory steered by it converges to a cycle $\mathcal{C}$, if the path $\mathcal{T}=\{\alpha(k):k=0,1,\cdots\}$ in $G$ keeps repeating $\mathcal{C}$ after a finite time.
+Define the weight and average weight of a path $\mathcal{T}=\{\delta_{N}^{a_{0}},\cdots,\delta_{N}^{a_{l}}\}\subseteq\mathcal{V}$, $l\geq 0$ as $w(\mathcal{T})=\sum_{i=0}^{l-1}w(\delta_{N}^{a_{i}},\delta_{N}^{a_{i+1}})$ and $\bar{w}(\mathcal{T})=w(\mathcal{T})/l$, respectively.
+Then, we have the following result.
+Theorem 5.
+MAS (11) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$, where $\mathcal{R}(\alpha_{0})$ denotes the set of constrained states reachable from $\alpha_{0}$.
+In addition, if MAS (11) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then
+  1. (i)
+    it holds $C(G[\Phi])\neq\emptyset$, where $\Phi=I(\Omega(s))\cap\mathcal{R}(\alpha_{0})$, $G[\Phi]=(\Phi,\mathcal{E}^{\prime},w)$ is the induced subgraph of $G$ induced by $\Phi$ with =$\mathcal{E}^{\prime}=\{(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}:\delta_{N}^{a},\delta_{N}^{b}\in\Phi\}$, and $C(\cdot)$ represents the set of all cycles in the graph;
+  2. (ii)
+    solutions to the infinite-horizon optimal control problem $P_{2}$ exist.